# SAR-Net: Multi-scale Direction-aware SAR Network via Global Information Fusion

### v1

@@ III-B2 Multi-level Fusion Module
Due to the typically high resolution and rich detailed information in SAR images, small objects manifest as shallow texture details, while large objects exhibit deep semantic information. Performing fusion only once cannot fully preserve the information of objects at different scales. Therefore, to enhance the model’s ability to detect multi-scale objects, we designed two multi-level fusion modules: the Shallow Multi-level Fusion Module (Shallow-MFM) and the Deep Multi-level Fusion Module (Deep-MFM). Shallow-MFM. Given that convolutional operations can effectively extract local features and texture information, which is crucial for detecting small objects in SAR images, Shallow-MFM utilizes Direction-aware Attention Block (DA Block). As shown in Fig. 4 (a), these blocks ensure the comprehensive capture and fusion of global detailed information in SAR images, followed by a separation operation. The formula is as follows:
@@
As shown in Fig. 2, for better control of the information propagation and fusion process, we only separate two features and compensate them onto the input features $C3$ and $C4$. This is because $C5$ contains the position and feature information of large objects, while the compensated features encompass more information about small objects. Simultaneously, the feature information from $C2$ has already been sufficiently utilized, and we no longer use it in the subsequent fusion process. Deep-MFM. Due to the advantages of transformers in handling global relationships and long-distance dependencies, and considering that large objects require more global information for accurate detection and localization, we employ a transformer fusion module in Deep-MFM. As illustrated in Fig. 4 (b), it consists of a stack of $N$ transformer blocks, where each transformer block includes Multi-Head Attention (MHA) and a Feed-Forward Network (FFN). We follow the configuration for Multi-Head Attention from [48], and this part of the process is expressed by the formula:
@@
here, $CB_{i}$ denotes the $i$-th convolution and batch normalization operation, $DW_{conv}$ represents Depth-Wise convolution, and $\psi$ is the ReLU activation function. Before merging the output features, to retain information about small objects in $K3$, we compensate the separated features onto the intermediate features $K4$ and $K5$, providing the detection head with output features rich in information about both small and large objects.
@@ III-C Direction-aware Attention Module
Research on object detection in SAR images has shown that attention mechanisms can significantly enhance model performance. However, existing attention mechanisms often concentrate solely on positional information, frequently neglecting directional cues. This results in a lack of a method that seamlessly integrates both directional and positional information and embeds them into the attention map. In response to this, we propose the direction-aware attention module. As shown in Fig. 6, it can be regarded as a means of enhancing the feature map’s representational capacity, obtaining direction-aware feature maps through two steps: Direction-aware Generation and Channel Attention Embedding. These steps contribute to facilitating the subsequent extraction of object features. Below, we provide a detailed description of these two steps.
@@ III-C1 Direction-aware Generation
Due to the dominance of background in SAR image content and the presence of interference areas such as docks, airports, and terminals, directly locating and detecting objects becomes challenging. Additionally, SAR images often exhibit the edge contours of objects, reflecting distinct directional features, while extracting detailed information within the objects is challenging. Therefore, leveraging the stability and distinctiveness of object direction information can effectively reduce the difficulty of localization. Based on the analysis above, deformable convolution, capable of adapting to and describing object deformations well, is employed. We perform deformable convolution along one direction (either the row or column direction) on the input features, generating two directional attention maps. These attention maps are then used to modulate the input features, and this process can be expressed as:
@@
here, $X$ represents the input features, and $\sigma$ denotes the Sigmoid function. This bidirectional convolution attention mechanism facilitates the extraction of directional information from low-level features in two directions, thereby guiding the refinement of input features. To further leverage the directional attention maps and avoid capturing only local contextual information, we forego using large-kernel convolutions and instead employ two parallel $1D$ feature encoding approaches. These methods independently aggregate input features in the vertical and horizontal directions, resulting in two separate localization-aware feature maps, as illustrated below:
@@ IV-A3 HRSID
To assess the ship detection capabilities of our proposed method across different datasets, we conducted performance validation on the HRSID dataset—a large-scale, high-resolution SAR image ship object detection dataset designed for tasks such as object detection and instance segmentation. The HRSID dataset encompasses SAR images with diverse properties, including variations in sea area, polarization, sea condition, resolution, and coastal port scenarios. As depicted in Fig. 7 (c), the dataset encompasses ships at multiple scales, annotated with bounding boxes within various environments (e.g., different polarization modes, sensor types, and scenes). Panoramic images with resolutions ranging from $1m$ to $5m$ were cropped to a uniform size of $800\times 800$. In total, the HRSID dataset consists of 5,604 cropped SAR images, containing 16,951 ships. These images are distributed in a ratio of 6.5:3.5 between the training and test sets. This dataset diversity ensures comprehensive evaluation and validation of our proposed method’s performance under various conditions and scenarios. Our proposed framework is implemented in PyTorch and runs on a workstation with an NVIDIA A100 GPU. During training, the batch size was set to 32, and the initial learning rate was 0.02, reduced to 0.0002 in a cosine decay manner, and trained for a total of 300 epochs to ensure it fully converged. Standard Stochastic Gradient Descent (SGD) was used as the optimizer, and the parameters were all at default values. For input images during training, we only used mosaics augment, discarding Hue Saturation Values (HSV) and random flipping as they are not beneficial for SAR images.
@@ IV-B Accuracy Metrics
TABLE VI: Comparison with Other Methods on SAR-AIRcraft-1.0. The Accuracy of Each Class is Represented by mAP with an Intersection Over the Union Threshold of 0.5
[…]
@@ IV-C1 Impact of Unity Compensation Mechanism
[…]
@@ IV-C3 Collaboration of the Framework
Figure 8: Visualization results of object detection under different methods on SAR-AIRcraft-1.0, (a)-(d) depict four randomly selected images, where differently colored rectangles represent aircraft of various types. To enhance clarity, we have boldened the object boxes. Yellow ellipses indicate false positives, while red ellipses represent instances of false negatives and misclassifications.
@@ IV-D Comparisons with Previous Methods
We used the SAR-AIRcraft-1.0 dataset to train and test several common benchmark detection methods, including two-stage detectors: Faster R-CNN [57], Cascade R-CNN [58], and one-stage detectors: YOLOv6 [47], YOLOv8 [59]. In addition to this, we also selected three latest SAR aircraft detection methods: SKG-Net [60], SA-Net [53] and MLSDNet [61], and a formidable model, HRLE-SARDet[62]. Table VI shows the detection results under different methods. In the Faster R-CNN method, the accuracy is 77.2%, which shows to a certain extent that the SAR-AIRcraft-1.0 dataset has a certain detection difficulty because aircraft objects of different categories have similar structures and sizes, and the same category also presents multi-scale characteristics. As shown in Table VI, it can be seen that due to the fusion of global information, SAR-Net achieves the highest performance in almost every category, and on this basis, we only add a small amount of computational overhead, but the performance has been greatly improved. Considering the advanced nature of YOLOv8, we also experimented with its variants from small to large versions. We observed that the performance fluctuated within a range of 0.3%, with a significant increase in computational cost. In addition, the detection accuracy of each category has certain differences. For example, A320/321 performs best in different algorithms compared to other categories. This is mainly due to the special size of A320/321, with a fuselage length of more than 40 meters, making it easy to distinguish. For some types of objects, such as ARJ21, due to its small size, the detailed features obtained are not sufficient, resulting in low detection accuracy.
@@ IV-D2 Comparison Results on SSDD
To verify the superiority of our method on different datasets, we selected the widely used SSDD dataset and compared it with some of the latest object detection methods and SAR ship detection methods. A variety of experiments ensured the rigor of our work. As indicated in the data presented in Table VII, our method exhibits strong competitiveness even when transferred to ship detection tasks. It can be observed that due to the relatively small size of the SSDD dataset, the current methods have already achieved a high level of average precision. However, our approach excels in information fusion and transmission, making it less prone to overfitting on small datasets. Consequently, this further elevates the benchmark performance of the dataset. Additionally, in comparison to the state-of-the-art (SOTA) model YOLOv8’s larger version, we experience only a marginal 0.1% decline in accuracy, while reducing parameters and GFLOPS to one-fourth and one-sixth, respectively.
@@ IV-D3 Comparison Results on HRSID
To further validate the generalization and robustness of our method, we opted for the more challenging ship detection dataset, HRSID, in comparison to SSDD. HRSID presents greater difficulty with a higher number of instances, larger image resolutions, and significant variations in object scales. This dataset serves as a more stringent test of the model’s capabilities in feature extraction and utilization. Continuing our evaluation, we chose state-of-the-art object detection methods, including those specifically tailored for ship detection in SAR images. Compared to other approaches, our method maintains excellent performance with a relatively small computational cost. Consistent with the results from SSDD, when compared to YOLOv8m, our method exhibits slightly lower accuracy but significantly smaller parameter size and GFLOPs, being one-fourth and one-sixth of YOLOv8m, respectively.
@@ IV-E Result Visualization
Due to the SAR-AIRcraft-1.0 dataset’s nature of fine-grained object detection, it presents certain detection challenges, leading to notable variations in detection results among different methods. To provide a more intuitive representation of the performance of various methods on SAR-AIRcraft-1.0, we selected a few images and visualized the detection results using different methods, as shown in Fig. 8. In comparison with other methods, our approach consistently identifies objects. While occasional false positives may occur, the incorporation of sufficient directional awareness information and the utilization of global fusion to minimize information loss result in minimal instances of false negatives in our detection results. For instance, the green bounding box (representing A220) in (b), the brown bounding box (representing ARJ21) in (c), and the orange bounding box (representing Boeing787) in (d) were undetected by other methods, highlighting the superiority of our approach in these scenarios. For the SSDD and HRSID datasets, given their high baseline levels of average precision and the single-category nature of object detection, we provide visualization results exclusively for SAR-Net. The visualizations for SSDD and HRSID are presented in Fig. 9 and Fig. 10, respectively. SAR-Net demonstrates excellent detection performance in both inshore and offshore backgrounds.

### v2

@@ III-B2 Multi-level Fusion Module
Due to the typically high resolution and rich detailed information in SAR images, small objects manifest as shallow texture details, while large objects exhibit deep semantic information. Performing fusion only once cannot fully preserve the information of objects at different scales. Therefore, to enhance the model’s ability to detect multi-scale objects, we designed two Multi-level Fusion modules: the Shallow Multi-level Fusion Module (Shallow-MFM) and the Deep Multi-level Fusion Module (Deep-MFM).
Shallow-MFM. Given that convolutional operations can effectively extract local features and texture information, which is crucial for detecting small objects in SAR images, Shallow-MFM utilizes Direction-aware Attention Block (DA Block). As shown in Fig. 4 (a), these blocks ensure the comprehensive capture and fusion of global detailed information in SAR images, followed by a separation operation. The formula is as follows:
@@
As shown in Fig. 2, for better control of the information propagation and fusion process, we only separate two features and compensate them onto the input features $C3$ and $C4$. This is because $C5$ contains the position and feature information of large objects, while the compensated features encompass more information about small objects. Simultaneously, the feature information from $C2$ has already been sufficiently utilized, and we no longer use it in the subsequent fusion process.
Deep-MFM. Due to the advantages of transformers in handling global relationships and long-distance dependencies, and considering that large objects require more global information for accurate detection and localization, we employ a transformer fusion module in Deep-MFM. As illustrated in Fig. 4 (b), it consists of a stack of $N$ transformer blocks, where each transformer block includes Multi-Head Attention (MHA) and a Feed-Forward Network (FFN). We follow the configuration for Multi-Head Attention from [48], and this part of the process is expressed by the formula:
@@
here, $CB_{i}$ denotes the $i$-th convolution and batch normalization operation, $DW_{conv}$ represents Depth-Wise convolution, and $\psi$ is the ReLU activation function.
Before merging the output features, to retain information about small objects in $K3$, we compensate the separated features onto the intermediate features $K4$ and $K5$, providing the detection head with output features rich in information about both small and large objects.
@@ III-C Direction-aware Attention Module
Research on object detection in SAR images has shown that attention mechanisms can significantly enhance model performance. However, existing attention mechanisms often concentrate solely on positional information, frequently neglecting directional cues. This results in a lack of a method that seamlessly integrates both directional and positional information and embeds them into the attention map. In response to this, we propose the direction-aware attention module.
As shown in Fig. 6, it can be regarded as a means of enhancing the feature map’s representational capacity, obtaining direction-aware feature maps through two steps: Direction-aware Generation and Channel Attention Embedding. These steps contribute to facilitating the subsequent extraction of object features. Below, we provide a detailed description of these two steps.
@@ III-C1 Direction-aware Generation
Due to the dominance of background in SAR image content and the presence of interference areas such as docks, airports, and terminals, directly locating and detecting objects becomes challenging. Additionally, SAR images often exhibit the edge contours of objects, reflecting distinct directional features, while extracting detailed information within the objects is challenging. Therefore, leveraging the stability and distinctiveness of object direction information can effectively reduce the difficulty of localization.
Based on the analysis above, deformable convolution, capable of adapting to and describing object deformations well, is employed. We perform deformable convolution along one direction (either the row or column direction) on the input features, generating two directional attention maps. These attention maps are then used to modulate the input features, and this process can be expressed as:
@@
here, $X$ represents the input features, and $\sigma$ denotes the Sigmoid function. This bidirectional convolution attention mechanism facilitates the extraction of directional information from low-level features in two directions, thereby guiding the refinement of input features.
To further leverage the directional attention maps and avoid capturing only local contextual information, we forego using large-kernel convolutions and instead employ two parallel $1D$ feature encoding approaches. These methods independently aggregate input features in the vertical and horizontal directions, resulting in two separate localization-aware feature maps, as illustrated below:
@@ IV-A3 HRSID
To assess the ship detection capabilities of our proposed method across different datasets, we conducted performance validation on the HRSID dataset—a large-scale, high-resolution SAR image ship object detection dataset designed for tasks such as object detection and instance segmentation. The HRSID dataset encompasses SAR images with diverse properties, including variations in sea area, polarization, sea condition, resolution, and coastal port scenarios. As depicted in Fig. 7 (c), the dataset encompasses ships at multiple scales, annotated with bounding boxes within various environments (e.g., different polarization modes, sensor types, and scenes). Panoramic images with resolutions ranging from $1m$ to $5m$ were cropped to a uniform size of $800\times 800$. In total, the HRSID dataset consists of 5,604 cropped SAR images, containing 16,951 ships. These images are distributed in a ratio of 6.5:3.5 between the training and test sets. This dataset diversity ensures comprehensive evaluation and validation of our proposed method’s performance under various conditions and scenarios.
Our proposed framework is implemented in PyTorch and runs on a workstation with an NVIDIA A100 GPU. During training, the batch size was set to 32, and the initial learning rate was 0.02, reduced to 0.0002 in a cosine decay manner, and trained for a total of 300 epochs to ensure it fully converged. Standard Stochastic Gradient Descent (SGD) was used as the optimizer, and the parameters were all at default values. For input images during training, we only used mosaics augment, discarding Hue Saturation Values (HSV) and random flipping as they are not beneficial for SAR images.
@@ IV-B Accuracy Metrics
TABLE VI: Comparison with Other Methods on SAR-AIRcraft-1.0. The Accuracy of Each category is Represented by mAP with an Intersection Over the Union Threshold of 0.5
Method A220 A320/321 A330 ARJ21 Boeing737 Boeing787 other mAP50(%) F1-score Params(M) GFLOPs
@@ IV-C1 Impact of Unity Compensation Mechanism
To study the effectiveness of each module in the unity compensation mechanism and verify the correctness of uniting deep and shallow layers’ global information, we conducted ablation experiments on three of the modules and measured the performance with mAP50 and F1-score, a common indicator in object detection. In addition, Parameter sizes and GFLOPs are also shown in Table II. We use YOLOv6n [47] as the baseline for ablation experiments to embed the modules of our unity compensation mechanism one by one. The feature alignment process is included in all experiments except the baseline. It can be seen that only when the Deep-MFM module is used, the accuracy is lower than the baseline. This is because it pays more attention to large objects, while most of the SAR images are small and medium-sized objects. In addition, other experiments have achieved performance improvements without significantly increasing the number of calculations, which proves the effectiveness of our global information fusion idea. In particular, when the MEM module is not used, we use simple addition to compensate for the original features.
@@ IV-C3 Collaboration of the Framework
Figure 8: Visualization results of object detection under different methods on SAR-AIRcraft-1.0, (a)-(d) depict four randomly selected images, where differently colored rectangles represent aircraft of various categories. To enhance clarity, we have boldened the object boxes. Yellow ellipses indicate false positives, while red ellipses represent instances of false negatives and misclassifications.
@@ IV-D Comparisons with Previous Methods
We used the SAR-AIRcraft-1.0 dataset to train and test several common benchmark detection methods, including two-stage detectors: Faster R-CNN [57], Cascade R-CNN [58], and one-stage detectors: YOLOv6 [47], YOLOv8 [59]. In addition to this, we also selected three latest SAR aircraft detection methods: SKG-Net [60], SA-Net [53] and MLSDNet [61], and a formidable model, HRLE-SARDet[62]. Table VI shows the detection results under different methods.
In the Faster R-CNN method, the accuracy is 77.2%, which shows to a certain extent that the SAR-AIRcraft-1.0 dataset has a certain detection difficulty because aircraft objects of different categories have similar structures and sizes, and the same category also presents multi-scale characteristics. As shown in Table VI, it can be seen that due to the fusion of global information, SAR-Net achieves the highest performance in almost every category, and on this basis, we only add a small amount of computational overhead, but the performance has been greatly improved. Considering the advanced nature of YOLOv8, we also experimented with its variants from small to large versions. We observed that the performance fluctuated within a range of 0.3%, with a significant increase in computational cost. In addition, the detection accuracy of each category has certain differences. For example, A320/321 performs best in different algorithms compared to other categories. This is mainly due to the special size of A320/321, with a fuselage length of more than 40 meters, making it easy to distinguish. For some objects in other categories, such as ARJ21 and Boeing aircraft, due to their multi-scale characteristics and varied directions, this pose a challenge for the network to learn invariant features within the same category. We enhance the detection capabilities of multi-scale and directional information-rich objects by fusing global information and performing compensation in deep and shallow layers respectively, as well as learning object direction-aware information.
@@ IV-D2 Comparison Results on SSDD
To verify the superiority of our method on different datasets, we selected the widely used SSDD dataset and compared it with some of the latest object detection methods and SAR ship detection methods. A variety of experiments ensured the rigor of our work.
As indicated in the data presented in Table VII, our method exhibits strong competitiveness even when transferred to ship detection tasks. It can be observed that due to the relatively small size of the SSDD dataset, the current methods have already achieved a high level of average precision. However, due to the continued presence of multi-scale objects and interference from background noise, this has prevented further performance improvements in other methods. Note that our method excels in information fusion and transmission, making it less prone to influence by these problems and less prone to overfit on small datasets. Consequently, this further elevates the benchmark performance of the dataset. Additionally, in comparison to the state-of-the-art (SOTA) model YOLOv8’s larger version, we still achieve a 0.4% precision advantage on mAP50, while reducing parameters and GFLOPS to one-fourth and one-sixth, respectively.
@@ IV-D3 Comparison Results on HRSID
To further validate the generalization and robustness of our method, we opted for the more challenging ship detection dataset, HRSID, in comparison to SSDD. HRSID presents greater difficulty with a higher number of instances, larger image resolutions, and significant variations in object scales. This dataset serves as a more stringent test of the model’s capabilities in feature extraction and utilization.
Continuing our evaluation, we chose state-of-the-art object detection methods, including those specifically tailored for ship detection in SAR images. Compared to other approaches, our method maintains excellent performance with a relatively small computational cost. Consistent with the results from SSDD, when compared to YOLOv8m, our method exhibits a 0.6% precision advantage on mAP50 and significantly smaller parameter size and GFLOPs, being one-fourth and one-sixth of YOLOv8m, respectively.
@@ IV-E Result Visualization
Due to the SAR-AIRcraft-1.0 dataset’s nature of fine-grained object detection, it presents certain detection challenges, leading to notable variations in detection results among different methods. To provide a more intuitive representation of the performance of various methods on SAR-AIRcraft-1.0, we selected a few images and visualized the detection results using different methods, as shown in Fig. 8. In comparison with other methods, our approach consistently identifies objects. While occasional false positives may occur, the incorporation of sufficient directional awareness information and the utilization of global fusion to minimize information loss result in minimal instances of false negatives in our detection results. For instance, the green bounding box (representing A220) in (b), the brown bounding box (representing ARJ21) in (c), and the orange bounding box (representing Boeing787) in (d) were undetected by other methods, highlighting the superiority of our approach in these scenarios.
For the SSDD and HRSID datasets, given their high baseline levels of average precision and the single-category nature of object detection, we provide visualization results exclusively for SAR-Net. The visualizations for SSDD and HRSID are presented in Fig. 9 and Fig. 10, respectively. SAR-Net demonstrates excellent detection performance in both inshore and offshore backgrounds.